OpenStudy (abbycross167):

Can someone please help me answer this question? I really need help... Solve the following problem using any method of solving systems. CHECK your solutions by substituting your coordinates into BOTH equations and simplifying. Question~ You want to plan a nutritious breakfast. It should supply at least 500 calories or more. Be sure your choices would provide a reasonable breakfast. Suppose you want to have cereal, milk, and one glass of tomato juice. Let c be the number of cups of cereal and m the number of cups of milk. The inequality 102c + 150m + 41 => 500 models the situation.

OpenStudy (abbycross167):

a) Determine three ordered pairs (c, m) that are solutions of the inequality where 0 <= c < 8 and 0 <= m < 4. b) Graph the inequality. (Assume you can eat partial eggs and drink partial apple juices.)

OpenStudy (abbycross167):

@jim_thompson5910 @retirEEd could one of y'all help me please?

jimthompson5910 (jim_thompson5910):

so you're stuck on part (a) ? or you figured that part out already?

OpenStudy (abbycross167):

Yes sir/ ma'am, part A is what I don't know what to do, part B I could figure out if I just get help with part A

jimthompson5910 (jim_thompson5910):

ok let's say c = 1. So that means you have 1 cup of cereal Plug in c = 1 and then solve the inequality for m. Tell me what you get

OpenStudy (abbycross167):

0 <= 1 < 8 is that the right way to replace c?

jimthompson5910 (jim_thompson5910):

I meant in \[\Large 102c + 150m + 41 \ge 500\]

OpenStudy (abbycross167):

Oh so.. 102(1)+150m+41≥500?

jimthompson5910 (jim_thompson5910):

yep, now solve for m

OpenStudy (abbycross167):

119/50?

jimthompson5910 (jim_thompson5910):

119/50 = 2.38 so you should find that m >= 2.38

jimthompson5910 (jim_thompson5910):

the restrictions on m are that `0 <= m < 4` so together we know `0 <= m < 4` and `m >= 2.38` you can pick any number for m that makes BOTH of the inequalities shown above to be true

jimthompson5910 (jim_thompson5910):

for example, you can pick m = 3 since it makes `m >= 2.38` true. Three is larger than 2.38 m = 3 also makes `0 <= m < 4` true because 3 is between 0 and 4

jimthompson5910 (jim_thompson5910):

There are infinitely many values of m to pick here

OpenStudy (abbycross167):

oh ok, so (3,4) would be a ordered pair for part A?

OpenStudy (abbycross167):

does 4 have to be one of the numbers in the ordered pairs?

jimthompson5910 (jim_thompson5910):

well recall that we made c = 1 so the first coordinate should be 1

jimthompson5910 (jim_thompson5910):

IF you picked m = 3, then one of the solutions would be (1,3)

OpenStudy (abbycross167):

oh ok... how would I get the other two ordered pairs?

jimthompson5910 (jim_thompson5910):

repeat the whole process over again but pick another value of c, say c = 2

jimthompson5910 (jim_thompson5910):

you can do this just as long as c is between 0 and 8 (excluding 8)

OpenStudy (abbycross167):

oh ok, and the other number can't be more than 4?

jimthompson5910 (jim_thompson5910):

or smaller than 0 based on the inequality 0 <= m < 4

OpenStudy (abbycross167):

102(2)+150m+41≥500 --- M>= 1.7 --- m=2 -- Ordered pair = (2,2) 102(5)+150m+41≥500 --- M>=0.34 --- m=1 -- ordered pair=(5,1)

OpenStudy (abbycross167):

@jim_thompson5910 is that the right way?

jimthompson5910 (jim_thompson5910):

let me check

OpenStudy (abbycross167):

ok

jimthompson5910 (jim_thompson5910):

I agree with the first one, but the second one I'm getting m >= -0.34 I would use c = 3 instead of c = 5

OpenStudy (abbycross167):

okay so, 102(3)+150m+41≥500 -- M>= 1.02 --- m= 3-- ordered pair = (3,3) ?

jimthompson5910 (jim_thompson5910):

much better

jimthompson5910 (jim_thompson5910):

so 3 solutions, of infinitely many, could be (1,3) (2,2) (3,3)

OpenStudy (abbycross167):

okay!! thank you sooo very much!! you're amazing!!

jimthompson5910 (jim_thompson5910):

you're welcome